# An Adversarial Non-Autoregressive Model for Text Generation with Incomplete Information

### v2

@@ 3.3 Dependency Feed Forward Network
Transformer (Vaswani et al., 2017) builds word dependencies solely based on the attention mechanism by assigning weights dynamically. This process, however, is unstable under the training of GANs. It will lead the models to lose word dependencies, and finally result in ungrammatical sentences. To tackle this problem, we propose Dependency Feed Forward Network (Dependency FFN) to strengthen the FFN module with the capacity of dependency modeling. The structure of Dependency FFN is shown in Figure 4 (b), and it is calculated as follows:
@@
where $\sigma(\cdot)$ is an activation function which is GELU in this work. With casual masks, $\mathbf{s}_{t-1}$ and $\mathbf{s}_{t}$ contain the information of first $(t-1)$ and $t$ words, respectively. Using the sum of these two variables can help the model to explicitly build stable dependencies between the $t$-th word and previous $(t-1)$ words in the fragile training process of GANs.
Table 1: FED and I. BLEU on the COCO Dataset and EMNLP Dataset.
Model DI COCO EMNLP
FED normal-$\downarrow$ I. BLEU normal-$\uparrow$ FED normal-$\downarrow$ I. BLEU normal-$\uparrow$
@@
In our experiment, adopting Dependency FFN in the discriminator can significantly improve model performance, while its effectiveness to the generator is limited. A powerful discriminator can guide the generator to model word dependencies with the original structure. We thus only adopt Dependency FFN in the discriminator.
@@
3.4 Extension to Conditional Generation
IIS not only covers unconditional generation tasks, but also includes conditional generation tasks. In the following, we introduce how to extend ANT to conditional generation. Given a condition representation $c$, the generator can consider it by shifting the original latent variable $z$. We find that using trainable factors to assign weights to $z$ and $c$ can slightly improve the performance. Thus, we incorporate the condition as follows:
$\hat{z}=\alpha_{1}\circ z+\alpha_{2}\circ c$ (6)
where $\alpha_{1}$ and $\alpha_{2}$ are two trainable variables. Then, $\hat{z}$ is fed into Position-Aware Self-Modulation, so the generator can consider the condition representation.
For the discriminator, we use the sum of word representations $x^{d}_{t}$ and conditional representations $c$ as the input: $\hat{x}^{d}_{t}=x^{d}_{t}+c$. Then, $\hat{x}^{d}_{t}$ is fed into the remaining modules of the discriminator.
4 Experiment
4.1 Experiment Setup
The experiment covers both unconditional generation and conditional generation to evaluate model performance under the IIS comprehensively. For the unconditional generation, we follow previous work (de Masson d’Autume et al., 2019; Ren and Li, 2022) and use sentences from two datasets: the COCO Image Caption Dataset (Lin et al., 2014) [^1] https://cocodataset.org and the EMNLP 2017 News Dataset [^2] http://www.statmt.org/wmt17/. The size of training sets of the COCO dataset and the EMNLP dataset are set to be 50,000 and 200,000, respectively. The COCO dataset can support evaluations in short sentence generation, while the EMNLP dataset focuses on long sentence generation. For the conditional generation, we randomly select 100,000 sentences from the Yelp Dataset [^3] https://www.yelp.com/dataset as training data and use emotion labels (positive or negative) as conditions.
@@ 4.1 Experiment Setup
4.2 Evaluation Metrics
The evaluation is conducted at both embedding level and token level. In embedding level, we use Universal Sentence Encoder [^4] https://tfhub.dev/google/universal-sentence-encoder/4 (Cer et al., 2018) to transform sentences into embeddings. Then, we calculate both Fréchet Embedding Distance (FED) (de Masson d’Autume et al., 2019) and Least Coverage Rate (LCR) (Ren and Li, 2022) to evaluate the overall similarity and the fine-grained similarity of two distributions, respectively.
In token level, we use Inverse-BLEU (I. BLEU) to evaluate model performance in terms of quality and diversity together. The Inverse-BLEU uses generated sentences as references and sentences in the test set as inferences. A model can get high Inverse BLEU only if it can obtain a good trade-off between quality and diversity (Ren and Li, 2022). We also draw a curve of BLEU (Papineni et al., 2002) and Self-BLEU (Zhu et al., 2018) by tuning the temperature of the model to evaluate the overall performance (Caccia et al., 2020). For conditional generation, we additionally use Universal Sentence Encoder to construct a classifier and use it to calculate Accuracy (Acc.), so as to verify whether the models generate sentences with expected labels or not.
4.3 Compared Model
Transformer (Vaswani et al., 2017) is an important compared model in our experiment, since it is now the mainstream model in various text generation tasks. Considering ANT is a GAN-based model, we also choose several representative models in language GANs for comparisons: RelGAN (Nie et al., 2019), which uses Gumbel-softmax to obtain gradients; ScratchGAN (de Masson d’Autume et al., 2019), which is based on REINFORCE; InitialGAN (Ren and Li, 2022), which does not use the above two method and adopts representation modeling. All the models mentioned-above are AR models whose Decoding Iteration (DI) is $O(N)$ ($N$ is the sequence length).
@@ 4.3 Compared Model
For the NAR models, we compare with NAGAN (Huang et al., 2021), which is also a GAN-based NAR model. Although existing MLE-based NAR models can not be adopted to the IIS directly, we still want to compare the differences between MLE-based methods and GAN-based methods. Thus, we use the idea of VAE to get the hidden representations of sentences and adapt two existing popular NAR models into the IIS. One is based on Non-autoregressive Transformer (NAT) (Gu et al., 2018) which uses the hidden representations to generate sentences in a fully NAR manner; we denote this model as V-NAT. Another one in based on conditional masked language model (CMLM) (Ghazvininejad et al., 2019) which generates sentences by iteratively masking and generating words; we denote this model as V-CMLM. The iteration number $k$ of V-CMLM is set to be 10 as in previous work (Ghazvininejad et al., 2019; Huang et al., 2022c). More illustrations about experiment details can be found in Appendix B. We will release our code to the public in the future.
4.4 Experimental Result
4.4.1 Unconditional Generation
The experimental results of the unconditional generation are shown in Table 1. “DI” indicates Decoding Iteration. For the COCO dataset, Transformer gets the best performance in AR models, while ANT is the best one in NAR models. ANT gets 0.013 in FED, which is better than other NAR models and close to AR models like ScratchGAN and InitialGAN. Similar results can be found in Inverse BLEU. ANT gets 31.12 in I. BLEU and it is much better than other NAR models. There are large gaps between MLE-based NAR models (V-NAT and V-CMLM) and AR models, and it is consistent with our theoretical analyses. The input with incomplete information increases the difficulties of MLE training. NAGAN, another GAN-based NAR model, is inferior to all the other models. It shows the limitations of the biased straight-through estimator.
For the EMNLP dataset, Transformer is still the best model. ANT outperforms other NAR models in FED, while V-CMLM can slightly outperform ANT in Inverse BLEU. The iterative decoding mechanism helps V-CMLM to better process complicated datasets with higher decoding latency. To further discuss their performance in the token level, we draw the curve of Self-BLEU and Negative BLEU by tuning the temperature in Transformer and show the results in Figure 5 (a). ANT is the only NAR model which can get comparable performance with AR models, while other NAR models (including V-CMLM) remain behind obviously. Specifically, NAGAN gets extremely low BLEU, which indicates that NAGAN can not generate fluency sentences. It reveals the difficulties of NAGAN to converge on complicated datasets. Furthermore, we compare Least Coverage Rate (LCR) of Transformer and other NAR models in Figure 5 (b). ANT outperform V-CMLM with lower decoding iterations, and it is the only NAR model which can get close performance with Transformer.
4.4.2 Conditional Generation
The experimental results of conditional generation are shown in Table 2. Transformer gets the best performance in all the evaluation metrics with more decoding iterations. Among NAR models, ANT gets comparable performance with V-CMLM in FED, and achieves higher Inverse BLEU and Accuracy with lower decoding latency. V-NAT, which has the same decoding iterations as ANT, is inferior to other models. Given an emotion label, there are a large number of possible candidates, and it will augment the inherent multi-modality problem in MLE-based NAR models. Thus, fully NAR models trained under MLE will easily mix words in different candidates and finally obtains poor performance. For the accuracy, ANT gets 88.35% which is the highest one among all the NAR models. ANT can generate sentences that are consistent with the given labels. It reveals that ANT has great potential in both unconditional generation and conditional generation.
@@ 4.4.3 Ablation Study
We design two features for ANT: 1) Dependency FFN, which strengthens model capacity in dependency modeling; and 2) Position-Aware Self-Modulation, which provides clearer signals for the model to recover the original data. We also conduct experiments to demonstrate the effectiveness of these two modules.
For Dependency FFN, we compare the performance between Dependency FFN and the original FFN on the COCO dataset and show the results in Figure 6 (a). ANT with Dependency FFN has lower FED and higher Inverse BLEU. It obtains better performance in both the token-level metrics and the embedding-level metrics. These results show that Dependency FFN can help improve model performance by modeling more accurate word dependencies.
@@
Figure 7: (a) Speedup of Different Models. (b) Case Study of Latent Interpolation
For Position-Aware Self-Modulation, we compare the training curves with original Self-Modulation with FED. The results are shown in Figure 6 (b). ANT with Position-Aware Self-Modulation converges faster, and finally achieves better performance. Position-Aware Self-Modulation can provide clearer input signals for the model and further enhance model performance.
4.5 Discussion
ANT is a fully NAR model so it has much lower decoding latency. We compare the speedup of different models and show the results in Figure 7 (a). ANT can generate sentences 14.75 times faster than Transformer. Even comparing with V-CMLM, it also has much lower decoding latency while obtaining comparable or even better performance.
Generative models can also boost the performance of classification models based on semi-supervised learning (SSL). We investigate the application of ANT in SSL by incorporating it into the training of a classification model. The classification model is trained to identify emotion labels of sentences in the Yelp dataset. We prepare two training sets. One is composed of 500 labeled data and the other one consists of 1,000 labeled data. The results are shown in Table 3. The classification models trained in SSL consistently outperform the ones trained in supervised learning (SL). ANT can obtain data following same distributions as the original data so as to help the classification model capture more accurate data distribution and achieve better performance.
Besides, ANT enables latent interpolation just like image GANs. There are two latent variables in ANT: $z$, which is sampled from a pre-defined distribution; and $c$, which is a condition representation. We fix one of them and gradually change the other one. The upper part of Figure 7 (b) shows the samples given by tuning $z$ with fixed $c$, in which ANT transforms one sentence into another one, with the middle sentences kept understandable. The lower part of Figure 7 (b) shows the samples given by changing $c$ from the negative representation to the positive representation. ANT gradually transforms negative words into positive ones while keeping the main structure of the sentence. Such latent interpolation is seldomly explored by NAR models, and it may inspire further ideas for related tasks.
@@ 5 Conclusion
In this work, we firstly analyze the limitations of existing NAR models in the IIS. The features of the IIS will augment the inherent multi-modality problems of MLE-based NAR models and the lower bounds between their learned distributions and real distributions will be higher under the IIS. Instead, we find GANs denote a more promising method to train NAR models in the IIS. Thus, we propose an Adversarial Non-autoregressive Transformer (ANT) based on GANs. ANT supports two novel features: Position-Aware Self-Modulation and Dependency FFN. With the help of these two facilities, ANT can get comparable performance as other AR models but with much lower decoding latency. Besides, we also demonstrate the great potential of ANT in various tasks like smooth latent interpolation and semi-supervised learning.
Although ANT can significantly reduce the decoding latency, it still does not outperform Transformer. In the future, we will explore more techniques to further improve the performance of ANT and extend it into other tasks.
Acknowledgments and Disclosure of Funding
The authors are grateful to Hongzhi Zhao for his help in mathematical proofs.
Appendix A Proof
Figure 8: Mapping Relations Described in Assumption 1.
Assumption 1 describes a relation which is shown in Figure 8. Giving an input, it has $n$ or $m$ possible mapping candidates and each candidate is consisted of $l$ elements. Based on it, we can obtain the following conditional probabilities:
$P_{data}(Y|X_{A})=P_{data}(y_{i}|X_{A})=1/n$ (7)
$P_{data}(Y|X_{B})=P_{data}(y_{i}|X_{B})=1/m$
It should be noted that $P_{data}(y_{i}|X_{A})$ and $P_{data}(y_{i}|X_{B})$ are not related to the sequence length $l$ in this scenario. Given a specific input, there are only $n$ or $m$ possible candidate elements at each position. Based on the conditional probabilities, we have:
$C_{A}$ $=\sum_{i=1}^{l}H_{data}(y_{i}|X_{A})-H_{data}(Y|X_{A})$ (8)
$=-l\cdot log(1/n)+log(1/n)$
$=-(l-1)\cdot log(1/n)$
$C_{B}$ $=\sum_{i=1}^{l}H_{data}(y_{i}|X_{B})-H_{data}(Y|X_{B})$ (9)
$=-l\cdot log(1/m)+log(1/m)$
$=-(l-1)\cdot log(1/m)$
Thus, we can calculate their difference as follows:
$C_{A}-C_{B}$ $=(l-1)\cdot[-log(1/n)+log(1/m)]$ (10)
$=(l-1)\cdot log(n/m)$
Thus, higher $l$ and $n$ lead to higher difference between $C_{A}$ and $C_{B}$. In addition to the theoretical analyses, our experimental results also demonstrate the difficulties of MLE-based NAR models in the IIS empirically.
Appendix B Experiment Details
B.1 Implementation Details
@@
The layer numbers of the aligner, generator and discriminator are all set to be 4. Their input dimension is 256, and the hidden dimension of FFN / Dependency FFN is 1,024. The head number is set to be 8. We use AdamW (Loshchilov and Hutter, 2019) as the optimizer of the aligner and the weight decay is set to be 1e-5; its learning rate is 0.0001. During the adversarial training, AdamW (Loshchilov and Hutter, 2019) is used as the optimizer of the discriminator whose weight decay is set to be 0.0001; its learning rate is 0.0002 for the COCO and Yelp dataset, and 0.00015 for the EMNLP dataset. We choose Adam (Kingma and Ba, 2015) as the optimizer of the generator and its learning rate is 0.0001. The $\beta_{1}$ and $\beta_{2}$ in the optimizers of the discriminator and the generator are set to be 0.5 and 0.9, respectively. The maximum training epoch is set to be 4,500. We implement our model based on Tensorflow [^5] https://www.tensorflow.org (Abadi et al., 2015) and the model is trained on NVIDIA GeForce RTX 3090.
B.2 V-NAT and V-CMLM

### v3

@@ 3.3 Dependency Feed Forward Network
Transformer (Vaswani et al., 2017) builds word dependencies by dynamically assigning weights in the attention mechanism. This process, however, is unstable under the training of GANs. It will lead the models to lose word dependencies, and finally result in ungrammatical sentences. We tackle this problem by proposing Dependency Feed Forward Network (Dependency FFN) to strengthen the FFN module with the capacity of dependency modeling. The structure of Dependency FFN is shown in Figure 3 (b), and calculated as follows:
@@
where $\sigma(\cdot)$ is an activation function which is GELU in this work. With causal masks, $\mathbf{s}_{t-1}$ and $\mathbf{s}_{t}$ contain the information of first $(t-1)$ and $t$ words, respectively. Using the sum of these two variables can help the model to explicitly build stable dependencies between the $t$-th word and previous $(t-1)$ words in the fragile training process of GANs.
3.4 Extension to Conditional Generation
Besides unconditional generation, conditional generation is frequently employed in a variety of tasks. We thus also extend ANT to conditional generation. Given a condition representation $c$, the generator can consider it by shifting the original latent variable $z$. We find that using trainable factors to assign weights to $z$ and $c$ can slightly improve the performance: $\hat{z}=\alpha_{1}\circ z+\alpha_{2}\circ c$, where $\alpha_{1}$ and $\alpha_{2}$ are two trainable variables. For the discriminator, we use the sum of word representations $x^{d}_{t}$ and conditional representations $c$ as the input: $\hat{x}^{d}_{t}=x^{d}_{t}+c$. Then, $\hat{x}^{d}_{t}$ is fed into the remaining modules of the discriminator.
4 Experiment
Table 1: FED and I. BLEU on the COCO Dataset and EMNLP Dataset (DI: Decoding Iteration).
Model DI COCO Dataset EMNLP Dataset
FED $\downarrow$ I. BLEU $\uparrow$ FED $\downarrow$ I.BLEU $\uparrow$
@@ 4 Experiment
4.1 Experiment Setup
The experiment covers both unconditional generation and conditional generation to evaluate model performance comprehensively. For the unconditional generation, the task is to generate sentences whose distribution can be as close as to the target sets. We follow the settings of previous work (de Masson d’Autume et al., 2019; Ren & Li, 2023) and use sentences from two datasets: the COCO Image Caption Dataset (Lin et al., 2014) [^1] https://cocodataset.org and the EMNLP 2017 News Dataset [^2] http://www.statmt.org/wmt17/. The size of training sets of the COCO dataset and the EMNLP dataset are set to be 50,000 and 200,000, respectively. The COCO dataset can support evaluations in short sentence generation, while the EMNLP dataset focuses on long sentence generation. For the conditional generation, we randomly select 100,000 sentences from the Yelp Dataset [^3] https://www.yelp.com/dataset as training data and use emotion labels (positive or negative) as conditions.
4.2 Evaluation Metrics
The evaluation is conducted at both embedding level and token level. In embedding level, we use Universal Sentence Encoder [^4] https://tfhub.dev/google/universal-sentence-encoder/4 (Cer et al., 2018) to transform sentences into embeddings. Then, we calculate both Fréchet Embedding Distance (FED) (de Masson d’Autume et al., 2019) and Least Coverage Rate (LCR) (Ren & Li, 2023) to evaluate the overall similarity and the fine-grained similarity of two distributions, respectively.
In token level, we use Inverse-BLEU (I. BLEU) to evaluate model performance in terms of quality and diversity together. Besides, we also draw a curve of BLEU (Papineni et al., 2002) and Self-BLEU (Zhu et al., 2018) by tuning the temperature of the model (Caccia et al., 2020). In the case of conditional generation, Accuracy (Acc.) is also employed to assess whether the models produce sentences that align with the input labels.
@@ 4.2 Evaluation Metrics
4.3 Compared Model
An important compared model is Transformer, which adopts AR structures and is trained on MLE. It is the mainstream model in various text generation tasks. Besides, a number of AR language GANs are also compared: SeqGAN (Yu et al., 2017), RankGAN (Lin et al., 2017), MaliGAN (Che et al., 2017), LeakGAN (Guo et al., 2018), and ScratchGAN (de Masson d’Autume et al., 2019), which are based on REINFORCE; RelGAN (Nie et al., 2019), which uses Gumbel-softmax to obtain gradients; InitialGAN (Ren & Li, 2023), which does not use the above two method and adopts representation modeling. All the models mentioned-above are AR models whose Decoding Iteration (DI) is $O(N)$ ($N$ is the sequence length).
For NAR models, we compare with another GAN-based model: NAGAN (Huang et al., 2021). Furthermore, two classical MLE-based NAR models are compared in our experiments: Non-autoregressive Transformer (NAT) (Gu et al., 2018) and conditional masked language model (CMLM) (Ghazvininejad et al., 2019). More illustrations about experiment details can be found in Appendix A. We will release our code to the public in the future.
4.4 Experimental Result
4.4.1 Unconditional Generation
The experimental results of the unconditional generation are shown in Table 1. For the COCO dataset, Transformer gets the best performance in AR models, while ANT is the best one in NAR models. More specifically, ANT obtains 0.013 in FED. This result outperforms a number of AR language GANs and is close to the InitalGAN, which is the best language GANs on the COCO dataset. Similar results can be found in Inverse BLEU (I. BELU). ANT gets 31.12 in I. BLEU and it is much better than other NAR models. The performance of MLE-based NAR models (NAT and CMLM) is far behind the AR models. Existing MLE-based NAR models make use of the characteristic of specific tasks (like strong corresponding relation between input and output in machine translation) to relieve the multi-modality problem Gu et al. (2018). Once they are transferred to more fundamental tasks (e.g., unconditional generation), the inhere problem will be more severe and lead to the obvious decrease of model performance. NAGAN, another GAN-based NAR model, is inferior to all the other models. It shows the limitations of the biased straight-through estimator.
For the EMNLP dataset, Transformer is still the best model. ANT outperforms other NAR models in FED, while CMLM can slightly outperform ANT in Inverse BLEU. The iterative decoding mechanism helps CMLM to better process complicated datasets with higher decoding latency. To further discuss their performance in the token level, we follow the suggestions from Caccia et al. (2020), and draw the curve of Self-BLEU and Negative BLEU by tuning the temperature in Transformer. The results are shown in Figure 4 (a). ANT is the only NAR model which can get comparable performance with AR models, while other NAR models (including CMLM) remain behind obviously. Specifically, NAGAN gets extremely low BLEU, which indicates that NAGAN cannot generate fluency sentences. It reveals the difficulties of NAGAN to converge on complicated datasets. Furthermore, we compare Least Coverage Rate (LCR) of Transformer and other NAR models in Figure 4 (b). ANT outperforms CMLM with lower decoding iterations, and it is the only NAR model which can get close performance with Transformer.
4.4.2 Conditional Generation
The experimental results of conditional generation are shown in Table 2. Among NAR models, ANT gets comparable performance with CMLM in FED, and achieves higher Inverse BLEU and Accuracy with lower decoding latency. NAT, which also generates samples in one decoding step, is inferior to other models. For the accuracy, ANT gets 88.35% which is the highest one among all the NAR models. ANT can generate sentences consistent with the given labels.
@@ 4.4.2 Conditional Generation
Both the experimental results in unconditional generation and conditional generation demonstrate the effectiveness of ANT. It outperforms a number of AR language GANs. Even comparing with the existing best language GANs, it can still obtain comparable performance in much lower latency. It is thus not necessary to build language GANs in AR structures. Besides, ANT is free from the theoretical limitations in MLE-based NAR models, so it can obtain better performance and denotes a more promising methods in building NAR models.
@@ 4.4.3 Ablation Study
We design two features for ANT: 1) Dependency FFN; and 2) Position-Aware Self-Modulation. Experiments are conducted to demonstrate the effectiveness of these two modules. For Dependency FFN, we compare the performance between Dependency FFN and the original FFN in Figure 5 (a). ANT with Dependency FFN has lower FED and higher Inverse BLEU. It obtains better performance in both the token-level metrics and the embedding-level metrics. These results show that Dependency FFN can help improve model performance by modeling more accurate word dependencies. For Position-Aware Self-Modulation, we compare the training curves with original Self-Modulation with FED in Figure 5 (b). ANT with Position-Aware Self-Modulation converges faster, and finally achieves better performance. Position-Aware Self-Modulation can enhance model performance by providing more diverse and effective representations.
4.5 Discussion
One advantage of ANT is that it only requires one decoding step and has high speedup. We compare the speedup of different models in Figure 6 (a). ANT is 14.75 times faster than Transformer. Even comparing with CMLM, it also has much lower decoding latency while obtaining comparable or even better performance. Besides, ANT has great potential in various applications. We explore the potential of ANT in different applications including semi-supervised learning and latent interpolation in the following.
@@ 4.5 Discussion
Generative models can be incorporated into semi-supervised learning (SSL) to assist the training of other models. It requires the models to be in high efficiency, since it needs to obtain new data during the training process. We investigate the application of ANT in SSL by incorporating it into the training of a classification model. The classification model is trained to identify emotion labels of sentences in the Yelp dataset. We prepare two training sets. One is composed of 500 labeled data and the other one consists of 1,000 labeled data. The results are shown in Table 3. For the classification model trained on 500 labeled data, its F1 score increases from 90.15% to 91.46%. It gains +1.31% improvement after using SSL. For the model trained 1,000 labeled data, its F1 scores increases from 91.87% to 93.62% in which +1.75% improvement is led by the SSL. The classification models trained in SSL consistently outperform the ones trained in supervised learning (SL). ANT can obtain data following same distributions as the original data so as to help the classification model improve performance.
Figure 6: (a) Speedup of Different Models. (b) Case Study of Latent Interpolation.
Besides, ANT enables latent interpolation just like image GANs. There are two latent variables in ANT: $z$, which is sampled from a pre-defined distribution; and $c$, which is a condition representation. We fix one of them and gradually change the other one. The first example in Figure 6 (b) shows the samples given by tuning $z$ with fixed $c$, in which ANT transforms one sentence into another one, with the middle sentences kept understandable. The second example shows the samples given by changing $c$ from the negative representation to the positive representation. ANT gradually transforms negative words into positive ones while keeping the main structure of the sentence. Such latent interpolation is seldomly explored by NAR models, and it may inspire further ideas for related tasks.
@@ 5 Conclusion
In this work, we firstly reveal the limitations in existing language GANs based on AR structures. The global optimality of GANs do not rely on specific structures and NAR structures can obtain similar performance with much lower latency. Then, we reveal two problems in building GAN-based NAR models: 1) Existing methods of integrating latent variables obtain similar representations which cannot describe the diversity of different words. 2) The attention mechanism in Transformer cannot build word dependencies stably. We tackle these two problems by proposing Position-Aware Self-Modulation and Dependency Feed Forward Network, respectively. Armed with these two facilities, we propose an Adversarial Non-autoregressive Transformer (ANT), a GAN-based NAR model. The experimental results demonstrate that ANT obtains comparable performance as other AR models but with much lower decoding latency. Besides, we also demonstrate the great potential of ANT in various tasks like smooth latent interpolation and semi-supervised learning.
Appendix A Experiment Details
A.1 Implementation Details
@@
The layer numbers of the aligner, generator and discriminator are all set to be 4. Their input dimension is 256, and the hidden dimension of FFN / Dependency FFN is 1,024. The head number is set to be 8. We use AdamW (Loshchilov & Hutter, 2019) as the optimizer of the aligner and the weight decay is set to be 1e-5; its learning rate is 0.0001. During the adversarial training, AdamW (Loshchilov & Hutter, 2019) is used as the optimizer of the discriminator whose weight decay is set to be 0.0001; its learning rate is 0.0002 for the COCO and Yelp dataset, and 0.00015 for the EMNLP dataset. We choose Adam (Kingma & Ba, 2015) as the optimizer of the generator and its learning rate is 0.0001. The $\beta_{1}$ and $\beta_{2}$ in the optimizers of the discriminator and the generator are set to be 0.5 and 0.9, respectively. The maximum training epoch is set to be 4,500. We implement our model based on Tensorflow [^5] https://www.tensorflow.org (Abadi et al., 2015) and the model is trained on NVIDIA GeForce RTX 3090.
A.2 Implementation of NAT and CMLM